Mathematics 243 – Algebraic Structures Definitions Quiz 1 – September 7, 2012

Your Name:

A) (5) What is the definition of the intersection of two sets: A ∩ B?

B) (5) Let f : {1, 2, 3} → {a, b, c, d, e} be defined by f(1) = b, f(2) = e and f(3) = b. What is the set f^{−1}({a, b, c})?